OpenStudy (anonymous):

Eliminate the parameter. x = t^2 + 2, y = t^2 - 4 I have no clue how to do this, thanks! Answer choices: y = x - 6, x ≥ 1 y = x + 6, x ≥ 1 y = x2 - 6, x ≥ 1 y = x2 + 6, x ≥ 1

OpenStudy (anonymous):

since \[x=t^{2}+2 \]\[y=t^{2}-4\] We can make t^{2} the subject which will give us\[x-2 = t^{2}\]\[y+4=t^{2}\]. Since t^{2} = t^{2} \[x-2=y+4\] Moving 4 to the left hand side to make y the subject will give us \[y=x-6\], hence the first option is correct.

OpenStudy (anonymous):

BTW we eliminated t^{2} because it was the parameter we were trying to get rid of...

OpenStudy (anonymous):

thanks so much! can you check this one too, i think i figured it out... Eliminate the parameter. x = 7t, y = t + 4 i got... y = x/7+ 4 @Mousam

OpenStudy (loser66):

yup

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

well done...

OpenStudy (anonymous):

okay thank you! @Loser66 @Mousam i have one more thats a little different and i cant quite figure it out... Eliminate the parameter. x = 5 cos t, y = 5 sin t

OpenStudy (anonymous):

make t the subject again

OpenStudy (anonymous):

so for x = 5cos(t), it will be t= cos^-1 (x/5)

OpenStudy (anonymous):

ok and then would y=sint would be t=sin^-1 (x/5)? @Mousam

OpenStudy (anonymous):

y=5sint*

OpenStudy (anonymous):

huh... if y=sin t then t= sin^-1 (y)

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

you've written t = sin^-1 (x/5)

OpenStudy (anonymous):

it should be sin^-1 (y/5)

OpenStudy (anonymous):

you there?

OpenStudy (anonymous):

ohh! ok yeah that makes since, ok so now i have cos^-1(x/5)=sin^-1 (y/5) right? @Mousam

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

but where do i go from there?

OpenStudy (anonymous):

since we know that t=cos^-1(x/5) and t=sin^1(y/5) cos^-1(x/5) = sin^-1(y/5) sin{cos^-1(x/5)} = y/5 5sin{cos^-1(x/5)} = y

OpenStudy (anonymous):

do you understand it?

OpenStudy (anonymous):

yes!! thanks SO much! the cos and sin were confusing me but i get it now, thanks for the help! @Mousam

OpenStudy (anonymous):

no worries... have fun and good luck

OpenStudy (loser66):

@Mousam to me, it's x^2 + y^2 = 25

OpenStudy (loser66):

because it forms the form of x = rcos theta, y = rsin theta.

OpenStudy (anonymous):

yea but all the examples he showed me didn't have any x^2 and y^2. i think they were just doing y=...x...

OpenStudy (anonymous):

so i just taught him that form...

OpenStudy (anonymous):

if @melper want the x^2, y^2 way then I have no problem teaching him the next time i meet him...